OpenStudy (anonymous):

MEDAL!!! What is the solution to sqrt5x+3=3sqrtx? A. -3/4 B. 3/4 C. -3/2 D. 3/2

OpenStudy (anonymous):

No its \[\sqrt{5x+3}=3\sqrt{x}\]